Direct Labor Variances Alpine Bicycle Company manufactures road bikes. The following data for September of the...

  1. Direct Labor Variances

    Alpine Bicycle Company manufactures road bikes. The following data for September of the current year are available:

    Quantity of direct labor used 480 hrs.
    Actual rate for direct labor $10.5 per hr.
    Bicycles completed in September 230
    Standard direct labor per bicycle 2 hrs.
    Standard rate for direct labor $10.7 per hr.

    a. Determine the direct labor rate variance and direct labor time variance. Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number.

    Rate variance $
    Time variance $
    Total direct labor cost variance $

    b. How much direct labor should be debited to Work in Process?
    $______
Answer #1

Rate Variance = (SR-AR) * AH

= (10.7-10.5) * 480

= - 96 Favourable (minus sign)

Time variance = (SH - AH) *SR

= (2*230 - 480) * 10. 7

= 214 Unfavorable

Total Cost variance = 96F+214U

= 118 Unfavorable

B.

Direct labor debited to wip

= (230*2*10.7)

= 4922

Comment if you face any issues